OpenStudy (anonymous):

Find the length of the diameter of a sphere with a volume of 767.15 ft3. 11.4 ft 5.7 ft 14.4 ft 7.5 ft

OpenStudy (anonymous):

The volume of a sphere is V = (4/3)(pi)(r^2) where r = radius, and V = volume. Plug in into the equation and find r, the riadius. The diameter is twice the radius.

OpenStudy (john_es):

You mean r^3.

OpenStudy (anonymous):

Sorry, yes, r^3. Sorry, my error in typing.

OpenStudy (anonymous):

V = (4/3) (pi)(r^3)

OpenStudy (john_es):

What @Easyaspi314 is true. This the first step. You can also find the value r, \[r=\sqrt[3]{\frac{3V}{4\pi}}\]And then, find the diameter, \[D=2r\]

OpenStudy (john_es):

So 11.4 is the correct answer.

Directrix (directrix):

767.15 =(4/3)* pi*r^3 r = 5.67 2r = 11.4 approx @John_ES is correct